# Parameter-Efficient Fine-Tuning via Circular Convolution

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) has gained popularity for fine-tuning large foundation models, leveraging low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ to represent weight changes (i.e., $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$). This method reduces trainable parameters and mitigates heavy memory consumption associated with full delta matrices by sequentially multiplying $\mathbf{A}$ and $\mathbf{B}$ with the activation. Despite its success, the intrinsic low-rank characteristic may limit its performance. Although several variants have been proposed to address this issue, they often overlook the crucial computational and memory efficiency brought by LoRA. In this paper, we propose Circular Convolution Adaptation (C3A), which not only achieves high-rank adaptation with enhanced performance but also excels in both computational power and memory utilization. Extensive experiments demonstrate that C3A consistently outperforms LoRA and its variants across various fine-tuning tasks.
@@ 1 Introduction
Figure 1: Relative comparison of C3A and baselines on RoBERTa-Base. The Pearson Correlation Coefficient (PCC) is evaluated on STS-B and the Matthew’s Correlation Coefficient (MCC) on CoLA. Accuracies across SST-2, MRPC, QNLI, and RTE are averaged and reported as Acc. #Param denotes the count of learnable parameters (excluding the head), while Memory indicates the peak GPU memory usage during training. Metrics highlighted in blue represent values where higher is better, and those in red indicate values where lower is better. Refer to Table 2 for absolute statistics.
In recent years, Large Foundation Models (LFMs) have surged in popularity due to their remarkable performance across various tasks in natural language processing (NLP) Brown et al. (2020); Touvron et al. (2023), computer vision (CV) Radford et al. (2021); Kirillov et al. (2023), and other domains Li et al. (2024). Characterized by their vast number of parameters and substantial computational demands, these models have set new benchmarks in accuracy and versatility. However, their size and complexity pose significant challenges for efficient fine-tuning, particularly in resource-constrained environments Malladi et al. (2023); Zhang et al. (2024b). To address this issue, Parameter-efficient fine-tuning (PEFT) techniques Mangrulkar et al. (2022), such as Low-Rank Adaptation (LoRA) Hu et al. (2021), have emerged as outstanding practical solutions.
LoRA reduces the number of trainable parameters by employing low-rank matrices to approximate weight changes, making fine-tuning more feasible and accessible without compromising the model’s performance. Concretely, LoRA can be expressed in math:
@@
where $\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$ are weight matrices, $\mathbf{B}\in\mathbb{R}^{d_{1}\times r},\mathbf{A}\in\mathbb{R}^{r\times d_{2}}$ are low-rank matrices to construct $\Delta\mathbf{W}$ and $\mathbf{x}\in\mathbb{R}^{d_{2}}$ is the activations. The number of trainable parameters is $r(d_{1}+d_{2})$, prompting the choice $r\ll\min(d_{1},d_{2})$ (e.g., $r=8$ for $d_{1}=d_{2}=1024$) for achieving high parameter efficiency. However, as discussed in Zeng & Lee (2023), LoRA’s capacity to represent a target model is strictly limited by $r$. To address this trade-off between performance and efficiency, Kopiczko et al. (2023) proposed Vector Random Matrix Adaptation (VeRA), which achieves competitive performance with significantly fewer trainable parameters through fixed random-matrix projections. Despite its low parameter count, VeRA requires substantial computational power and memory due to the large size of the random matrices used for projection. This raises the following open research question in the context of PEFT:
@@
How can high-rank adaptation be achieved without incurring significant costs in terms of time complexity, memory usage, and trainable parameters?
As discussed in Dosovitskiy et al. (2020), dense linear layers lack inductive biases, which poses challenge on the optimization process and therefore limits transformers’ ability to compete with Convolutional Neural Networks (CNNs) when data is scarce. Given the constraints of a small training set for the downstream task, we posit that incorporating a stronger inductive bias could improve the adaptation performance. In light of this, we propose Circular Convolution Adaptation (C3A), featuring the circular convolution operator Bamieh (2018).
Figure 2: Overview of LoRA (A) and our C3A (B,C) method. In LoRA, only low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ are trained and the delta weight is represented by their product (i.e., $\Delta\mathbf{W}=\mathbf{BA}$). The total trainable parameter number is $r(d_{1}+d_{2})$, which is assosiated with the rank of the delta weight. In C3A, circular convolution kernels $\Delta\mathbf{w}$ are tuned to adapt to the downstream task and the delta weight is represented by the (block-)circular matrix they construct (i.e., $\Delta\mathbf{W}=\mathcal{C}_{\mathrm{(blk)}}(\Delta\mathbf{w})$). The total trainable parameter count is $\frac{d_{1}d_{2}}{b}$, which disentangles with the rank of the delta weight. Here, $b$ is the block size of the block-circular matrix and it should be a common divisor (CD) of $d_{1}$ and $d_{2}$.
Circular convolution has been famous in signal processing Li et al. (2020) and cryptography Dworkin et al. (2001) for its distinguished efficiency and compactness. This operator can be equivalently represented as multiplication by a circulant matrix, which offers rank flexibility that is unconstrained by the number of trainable parameters. Moreover, leveraging the Fast Fourier Transform (FFT), C3A ensures better time and memory efficiency than directly multiplying the circulant matrix Bamieh (2018), rendering it competitive with LoRA from the efficiency aspect. In a word, circular convolution is a promising option for escaping the rank curse of LoRA at minimal cost. We summarize our contribution as follows:
@@
① We introduce C3A, a novel approach for PEFT using convolution. This method leverages the circular convolution operation and its equivalent circulant matrix to provide a flexible rank, which is free of linear constraint by the number of trainable parameters, for the delta matrix.
@@
② Leveraging the elegant diagonalization of the circulant matrix, we implement both the forward pass and backpropagation using FFT. With the incorporation of FFT, the computation and memory efficiency of C3A excels. C3A strikes a unique balance between performance an efficiency.
@@
③ To offer greater flexibility in controlling the number of trainable parameters, we extend C3A by incorporating block-circular convolution, which results in block-circulant matrices. This extension allows C3A to achieve fully customizable parameter counts as well as adaptable rank configurations.
@@
④ We validate C3A through comprehensive fine-tuning experiments across diverse tasks including natural language understanding, instruction tuning and image classification. Experiments demonstrate C3A’s outstanding accuracy and memory merits compared to existing state-of-the-art methods.
@@ 2.1 Parameter-Efficient Fine-Tuning
Research on Parameter-Efficient Fine-Tuning (PEFT) has generally progressed along three main directions. The first direction involves partially updating the pre-trained neural network (e.g., the body Oh et al. (2020) or the biases Zaken et al. (2021)). Traditional methods relied on hand-crafted heuristics Han et al. (2015); Oh et al. (2020) to identify which parameters are crucial and should be fine-tuned. More advanced approaches employ optimization techniques Guo et al. (2020); Xu et al. (2021); Fu et al. (2023). For example, Guo et al. (2020) reformulated such a discrete optimization problem into a continuous one by employing Bernoulli masks and the Gumbel-softmax approximation.
The second direction emerged to maintain the integrity of the pre-trained model while enabling a high degree of parameter sharing through adapter-based methods He et al. (2021); Rebuffi et al. (2017); Rücklé et al. (2020); Liu et al. (2022). These works focus on integrating additional modules, termed adapters, to fit the downstream task, effectively decoupling the pre-trained model parameters from those specific to the downstream task. Prompt Tuning Brown et al. (2020); Gao et al. (2020); Chen et al. (2023); Zhang et al. (2024a) and Prefix Tuning Li & Liang (2021) also fall into this category, despite ignoring potential semantic meanings.
The final direction is characterized by delta-weight-based methods, such as Low-Rank Adaptation (LoRA) Hu et al. (2021) and Orthogonal Fine-tuning (OFT) Qiu et al. (2023). These methods bridge the gap between the pre-trained model and the downstream task by adaptive delta weights, which are stored separately while used in combination with the pre-trained weights. This unique design enables disentanglement of the pretrained and downstream-specific weights. Namely, it achieves parameter sharing and preserves the ability to integrate without additional inference cost. LoRA models the delta-weights by an additive matrix while OFT does it by a multiplicative one. To further improve either parameter efficiency or performance, many variants has been proposed for both of the methods Kopiczko et al. (2023); Liu et al. (2024; 2023); Yuan et al. (2024). However, these methods can hardly achieve high parameter efficiency and performance without incurring heavy computation and memory usage.
@@ 2.2 Circular Convolution
Circular convolution has been extensively studied in signal processing Rabiner et al. (1978); McGillem & Cooper (1984); Li et al. (2020) and cryptography Dworkin et al. (2001); Gong et al. (2024). Owing to its computational advantages, circular convolution has also been explored in machine learning for generating long embeddings of high-dimensional data Yu et al. (2014) and compressing heavily parameterized layers Cheng et al. (2015); Ding et al. (2017). Remarkably, it achieves these efficiencies without significant performance degradation, which makes it a promising technique for fine-tuning applications.
Despite its success in small neural networks such as LeNet Cheng et al. (2015), circular convolution has not demonstrated lossless performance in modern large foundational models (LFMs) or even in their base architecture, the transformer. This limitation may be attributed to the conflict between its high intrinsic bias (i.e., the circulant pattern) and the vast amount of data required for training LFMs. Conversely, when fine-tuning LFMs, it is often impractical to collect as much data as needed for training from scratch. In such scenarios, the intrinsic bias of circular convolution could potentially serve as a regularization mechanism, thereby benefiting the optimization process of fine-tuning.
@@ 3 Method
In this section, we present C3A (see an overview in Figure 2), a novel PEFT approach based on the circular convolution. C3A follows LoRA’s setting of learning an additive linear operation over the original dense linear transformation. However, instead of using low-rank decomposition and the matrix multiplication operator, C3A resorts to circular convolution as this additive linear operation. Section 3.1 introduces the notations we use. Section 3.2 discusses the circular convolution operator, its equivalent circulant matrix, and its calculation in the frequency domain. Section 3.3 details an efficient method for backpropagation. Section 3.4 describes block-wise convolution for controlling the number of trainable parameters. Finally, Section 3.5 analyzes the computational complexity.
@@ 3.1 Notations
The adapted weight matrix, the original weight matrix, and the delta matrix are denoted by $\mathbf{W}$, $\mathbf{W}_{0}$, and $\Delta\mathbf{W}$, respectively ($\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$). The activation vector of the previous layer is denoted by $\mathbf{x}\in\mathbb{R}^{d_{2}}$. The post-transformation vector is $\mathbf{z}$, where $\mathbf{z}=\mathbf{W}\mathbf{x}\in\mathbb{R}^{d_{1}}$, and the incremental part is denoted by $\Delta\mathbf{z}$, where $\Delta\mathbf{z}=\Delta\mathbf{W}\mathbf{x}\in\mathbb{R}^{d_{1}}$. The matrices $\mathbf{A}$ and $\mathbf{B}$ are low-rank matrices introduced by LoRA to represent $\Delta\mathbf{W}$, with $r$ being their rank. $r_{v}$ specifies the rank of the random projection matrix used in VeRA. The circular convolution kernel of C3A is denoted by $\Delta\mathbf{w}$ and the circular convolution operator by $\star$. The loss function is represented by $\mathcal{L}$. The Fast Fourier Transform and its inverse are denoted by $\mathrm{FFT}$ and $\mathrm{iFFT}$, respectively. The Hadamard product is denoted by $\circ$.
@@ 3.2 Circular Convolution
Theoretically, the rank of $\mathcal{C}(\Delta\mathbf{w})$ is given by $d-\mathrm{Deg}(\mathrm{gcd}(f(x),x^{d}-1))$ Ingleton (1956), where $\mathrm{Deg}(\cdot)$ denotes the degree of a polynomial, $f(x)$ is the polynomial associated with $\Delta\mathbf{w}$ (i.e., $f(x)=\sum_{i=1}^{d}\Delta w_{i}x^{i-1}$), and $\mathrm{gcd}(\cdot)$ represents the greatest common divisor. Consequently, the theoretical upper bound on the rank of $\mathcal{C}(\Delta\mathbf{w})$ is $d$. By learning $\Delta\mathbf{w}$ in the $\mathbb{R}^{n}$ oracle, C3A automatically achieves dynamic rank selection, which is not linearly constrained by the number of learnable parameters, unlike LoRA.
To achieve high efficiency, we leverage the beautiful circulant structure of $\mathcal{C}(\Delta\mathbf{w})$, which makes it diagonalizable by the Fourier basis ($\mathbf{F}$) . In math, it can be described as $\mathcal{C}(\Delta\mathbf{w})=\mathbf{F}\frac{\Lambda}{d}\mathbf{F}^{-1}$ Golub & Van Loan (1996), where $\Lambda$ is its eigenvalues and can be calculated by a Fourier transform of the first row (i.e., $\Lambda=\mathrm{diag}(\mathbf{F}\Delta\mathbf{w})$). Therefore, we can calculate $\Delta\mathbf{w}\star\mathbf{x}$ as

### v3

@@ Abstract
Low-Rank Adaptation (LoRA) has gained popularity for fine-tuning large foundation models, leveraging low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ to represent weight changes (i.e., $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$). This method reduces trainable parameters and mitigates heavy memory consumption associated with full delta matrices by sequentially multiplying $\mathbf{A}$ and $\mathbf{B}$ with the activation. Despite its success, the intrinsic low-rank characteristic may limit its performance. Although several variants have been proposed to address this issue, they often overlook the crucial computational and memory efficiency brought by LoRA. In this paper, we propose Circular Convolution Adaptation (C3A), which not only achieves high-rank adaptation with enhanced performance but also excels in both computational power and memory utilization. Extensive experiments demonstrate that C3A consistently outperforms LoRA and its variants across various fine-tuning tasks. Code released.
Parameter-Efficient Fine-Tuning via Circular Convolution
Aochuan Chen[Uncaptioned image]thanks: Equal contribution. Jiashun Cheng[Uncaptioned image]11footnotemark: 1 Zijing Liu[Uncaptioned image] Ziqi Gao[Uncaptioned image] Fugee Tsung[Uncaptioned image] Yu Li[Uncaptioned image] Jia Li[Uncaptioned image]thanks: Corresponding author. [Uncaptioned image]The Hong Kong University of Science and Technology (Guangzhou) The Hong Kong University of Science and Technology [Uncaptioned image]International Digital Economy Academy jialee@hkust-gz.edu.cn
@@ 1 Introduction
Figure 1: Performance comparison of C3A and other methods relative to LoRA on LLaMA-8B. Higher values are better across all metrics. See Table 3 and Table 4 for more statistics.
In recent years, Large Foundation Models (LFMs) have witnessed a pronounced ascendance in both scholarly and practical realms, attributable to their exceptional efficacy across diverse tasks in natural language processing (NLP) (Brown et al., 2020; Touvron et al., 2023), computer vision (CV) (Radford et al., 2021; Kirillov et al., 2023), and other domains (Li et al., 2024a, b, 2025). Distinguished by an extensive parameter count and significant computational requisites, these models have established unprecedented benchmarks in both accuracy and versatility. Nonetheless, their considerable size and intricate structure present formidable obstacles for efficient fine-tuning, especially within resource-constrained environments (Malladi et al., 2023; Zhang et al., 2024b). To mitigate these challenges, parameter-efficient fine-tuning (PEFT) techniques (Mangrulkar et al., 2022), exemplified by Low-Rank Adaptation (LoRA) (Hu et al., 2021), have emerged as highly effective solutions.
LoRA reduces the number of trainable parameters by leveraging low-rank matrices to approximate alterations in weights, thereby facilitating fine-tuning without degrading the model’s efficacy. Specifically, LoRA can be articulated mathematically as follows:
@@
where $\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$ are weight matrices, $\mathbf{B}\in\mathbb{R}^{d_{1}\times r},\mathbf{A}\in\mathbb{R}^{r\times d_{2}}$ are low-rank matrices formulated to construct $\Delta\mathbf{W}$, and $\mathbf{x}\in\mathbb{R}^{d_{2}}$ are the activations. The number of trainable parameters is $r(d_{1}+d_{2})$, thereby motivating the selection of $r\ll\min(d_{1},d_{2})$ (e.g., $r=8$ for $d_{1}=d_{2}=1024$) to attain elevated parameter efficiency. Nonetheless, as elaborated by Zeng and Lee (2023), the potential of LoRA to encapsulate a target model is inherently constrained by $r$. In an effort to reconcile the dichotomy between performance and efficiency, Kopiczko et al. (2023) introduced Vector Random Matrix Adaptation (VeRA). VeRA attains comparable performance with a markedly reduced count of trainable parameters via fixed random-matrix projections. However, despite its minimal parameter count, VeRA demands considerable computational resources and memory capacity due to the extensive nature of the random matrices employed for projection. As depicted in Figure 1, other representative works share the same resource problem. This precipitates the following open research question within the scope of PEFT:
@@
Beyond low parameter counts, how to achieve high-rank adaptation without incurring significant costs of time and memory?
To address this question, we introduce Circular Convolution Adaptation (C3A), which incorporates the circular convolution operator (Bamieh, 2018). Circular convolution has garnered significant attention in both signal processing (Li et al., 2020) and cryptography (Dworkin et al., 2001) due to its exceptional efficiency and compactness. This operator can be equivalently expressed as multiplication by a circulant matrix, providing rank flexibility that is independent of the number of trainable parameters. Furthermore, by employing the Fast Fourier Transform (FFT), C3A achieves superior time and memory efficiency compared to the direct multiplication (Bamieh, 2018), which makes it competitive with LoRA in terms of efficiency.
In addition, as explicated by Dosovitskiy et al. (2020), dense linear layers exhibit a deficiency of inductive biases, engendering a complex optimization landscape. Consequently, this hampers the effectiveness of transformers in comparison to Convolutional Neural Networks (CNNs) under conditions of limited data availability. Within the framework of a constrained training dataset for the downstream task, we postulate that a robust inductive bias could potentially augment adaption performance. The circular pattern in C3A serves precisely as such an inductive bias.
In summary, circular convolution presents a promising solution for circumventing the rank limitations of LoRA at minimal costs. Our contributions can be summarized as follows:
@@
❶ We introduce C3A, a novel approach for PEFT. This method leverages the circular convolution operation and its equivalent circulant matrix to provide a flexible rank, which is free of linear constraint by the number of trainable parameters.
@@
❷ Leveraging the elegant diagonalization of the circulant matrix, we implement both the forward pass and backpropagation using FFT. With the incorporation of FFT, the computation and memory efficiency of C3A excels. C3A strikes a unique balance between performance an efficiency.
@@
❸ To offer greater flexibility in controlling the number of trainable parameters, we extend C3A by incorporating block-circular convolution, which results in block-circulant matrices. This extension allows C3A to achieve fully customizable parameter counts as well as adaptable rank configurations.
@@
❹ We validate C3A through comprehensive fine-tuning experiments across diverse tasks, which demonstrate C3A’s outstanding accuracy and memory merits compared to existing methods.
@@ 2.1 Parameter-Efficient Fine-Tuning
Research on PEFT has generally progressed along three main directions. The first direction involves partially updating the pre-trained neural network (e.g., the layer norm (Basu et al., 2024) or the biases (Zaken et al., 2021)). Traditional methods relied on hand-crafted heuristics (Raghu et al., 2019) to identify which parameters are crucial and should be fine-tuned. More advanced approaches employ optimization techniques (Guo et al., 2020; Xu et al., 2021; Fu et al., 2023). For example, Guo et al. (2020) reformulated such a discrete optimization problem into a continuous one by employing Bernoulli masks and the Gumbel-softmax approximation (Jang et al., 2016).
The second direction emerged to maintain the integrity of the pre-trained model while enabling a high degree of parameter sharing through adapter-based methods (He et al., 2021; Rebuffi et al., 2017; Rücklé et al., 2020; Liu et al., 2022; Lian et al., 2022). These works focus on integrating additional modules, termed adapters, to fit the downstream task, effectively decoupling the pre-trained model parameters from those specific to the downstream task. Prompt Tuning (Brown et al., 2020; Gao et al., 2020; Chen et al., 2023; Zhang et al., 2024a) and Prefix Tuning (Li and Liang, 2021; Jia et al., 2022) also fall into this category, despite ignoring potential semantic meanings.
The final direction is characterized by delta-weight-based methods, such as Low-Rank Adaptation (LoRA) (Hu et al., 2021) and Orthogonal Fine-tuning (OFT) (Qiu et al., 2023). These methods bridge the gap between the pre-trained model and the downstream task by adaptive delta weights, which are stored separately while used in combination with the pre-trained weights. This unique design enables disentanglement of the pretrained and downstream-specific weights. Namely, it achieves parameter sharing and preserves the ability to integrate without additional inference cost. LoRA models the delta-weights by an additive matrix while OFT does it by a multiplicative one. To further improve either parameter efficiency or performance, many variants has been proposed for both of the methods (Kopiczko et al., 2023; Liu et al., 2024b, 2023; Yuan et al., 2024; Hayou et al., 2024b). However, these methods can hardly achieve high parameter efficiency and performance without incurring heavy computation and memory usage.
@@ 2.2 Circular Convolution
Circular convolution has been extensively studied in signal processing (Rabiner et al., 1978; McGillem and Cooper, 1984; Li et al., 2020) and cryptography (Dworkin et al., 2001; Gong et al., 2024). Owing to its computational advantages, circular convolution has also been explored in machine learning for generating long embeddings of high-dimensional data (Yu et al., 2014) and compressing heavily parameterized layers (Cheng et al., 2015; Ding et al., 2017). Remarkably, it achieves these efficiencies without significant performance degradation, which makes it a promising technique for fine-tuning applications.
Despite its success in small neural networks such as LeNet (Cheng et al., 2015), circular convolution has not demonstrated lossless performance in modern large foundational models (LFMs) or even in their base architecture, the transformer. This limitation may be attributed to the conflict between its high intrinsic bias (i.e., the circulant pattern) and the vast amount of data required for training LFMs. Conversely, when fine-tuning LFMs, it is often impractical to collect as much data as needed for training from scratch. In such scenarios, the intrinsic bias of circular convolution could potentially serve as a regularization mechanism, thereby benefiting the optimization process of fine-tuning.
@@ 3 Method
Figure 2: Overview of LoRA (A) and our C3A (B,C) method. In LoRA, only low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ are trained and the delta weight is represented by their product (i.e., $\Delta\mathbf{W}=\mathbf{BA}$). The total trainable parameter number is $r(d_{1}+d_{2})$, which is assosiated with the rank of the delta weight. In C3A, circular convolution kernels $\Delta\mathbf{w}$ are tuned to adapt to the downstream task and the delta weight is represented by the (block-)circular matrix they construct (i.e., $\Delta\mathbf{W}=\mathcal{C}_{\mathrm{(blk)}}(\Delta\mathbf{w})$). The total trainable parameter count is $\frac{d_{1}d_{2}}{b}$, which disentangles with the rank of the delta weight. Here, $b$ is the block size and it should be a common divisor (CD) of $d_{1}$ and $d_{2}$.
In this section, we present C3A (see an overview in Figure 2), a novel PEFT approach based on the circular convolution. C3A follows LoRA’s setting of learning an additive linear operation over the original dense linear transformation. However, instead of using low-rank decomposition and the matrix multiplication operator, C3A resorts to circular convolution as this additive linear operation.
@@ 3.1 Notations
The adapted weight matrix, the original weight matrix, and the delta matrix are denoted by $\mathbf{W}$, $\mathbf{W}_{0}$, and $\Delta\mathbf{W}$, respectively ($\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$). The activation vector of the previous layer is denoted by $\mathbf{x}\in\mathbb{R}^{d_{2}}$. The post-transformation vector is $\mathbf{z}$, where $\mathbf{z}=\mathbf{W}\mathbf{x}\in\mathbb{R}^{d_{1}}$, and the incremental part is denoted by $\Delta\mathbf{z}$, where $\Delta\mathbf{z}=\Delta\mathbf{W}\mathbf{x}\in\mathbb{R}^{d_{1}}$. The matrices $\mathbf{A}$ and $\mathbf{B}$ are low-rank matrices introduced by LoRA to represent $\Delta\mathbf{W}$, with $r$ being their rank. $r_{v}$ specifies the rank of the random projection matrix used in VeRA. The circular convolution kernel of C3A is denoted by $\Delta\mathbf{w}$ and the circular convolution operator by $\star$. The loss function is represented by $\mathcal{L}$. The Fast Fourier Transform and its inverse are denoted by $\mathrm{FFT}$ and $\mathrm{iFFT}$. The Hadamard product is denoted by $\circ$.
@@ 3.2 Circular Convolution
Theoretically, the rank of $\mathcal{C}(\Delta\mathbf{w})$ is given by $d-\mathrm{Deg}(\mathrm{gcd}(f(x),x^{d}-1))$ (Ingleton, 1956), where $\mathrm{Deg}(\cdot)$ denotes the degree of a polynomial, $f(x)$ is the polynomial associated with $\Delta\mathbf{w}$ (i.e., $f(x)=\sum_{i=1}^{d}\Delta w_{i}x^{i-1}$), and $\mathrm{gcd}(\cdot)$ represents the greatest common divisor. Consequently, the theoretical upper bound on the rank of $\mathcal{C}(\Delta\mathbf{w})$ is $d$. By learning $\Delta\mathbf{w}$ in the $\mathbb{R}^{n}$ oracle, C3A automatically achieves dynamic rank selection, which is not linearly constrained by the number of learnable parameters, unlike LoRA.
To achieve high efficiency, enlightened by Ding et al. (2017), we leverage the beautiful circulant structure of $\mathcal{C}(\Delta\mathbf{w})$, which makes it diagonalizable by the Fourier basis ($\mathbf{F}$) . In math, it can be described as $\mathcal{C}(\Delta\mathbf{w})=\mathbf{F}\frac{\Lambda}{d}\mathbf{F}^{-1}$ (Golub and Van Loan, 1996), where $\Lambda$ is its eigenvalues and can be calculated by a Fourier transform of the first row (i.e., $\Lambda=\mathrm{diag}(\mathbf{F}\Delta\mathbf{w})$). Therefore, we can calculate $\Delta\mathbf{w}\star\mathbf{x}$ as